OpenStudy (anonymous):

Simplify f(x) = -6x + 11 f(-3/2) g(x) = x^2 - 5x - 3 g(-2)

OpenStudy (anonymous):

Are these 2 different questions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(-3/2)=-6(-3/2)+11=9+11=20 g(-2)=(-2)^2-5(-2)-3=4+10-3=11 @AlysaMckee

OpenStudy (anonymous):

\[f(\frac{ -3 }{ 2 }) = -6(\frac{ -3 }{ 2) }+11\] \[f(\frac{ -3 }{ 2 }) = -3(-3) + 11\] \[f(\frac{ -3 }{ 2 }) = 9 + 11 =20\]